OpenStudy (anonymous):

y=-x^2-6x how do I find the vertex?

OpenStudy (campbell_st):

the vertex lies on the line of symmetry. the line of symmetry is found using \[x = \frac{-b}{2a}\] in your question a = -1 and b = -6 so the line of symmetry is \[x = \frac{- -6}{2 \times (-1)} = -3\] this is the x value in the vertex substitute x = -3 into the equation to find the y=value of the vertex

OpenStudy (anonymous):

i came up with 9

OpenStudy (helder_edwin):

it easier to complete the square (i don't know if you this expresion, it is a literal translation from spanish), but anyway \[ y=-x^2-6x \] \[ y-9=-(x^2+6x+9) \] \[ y-9=-(x+3)^2 \] now from this, what is the vertex?

OpenStudy (campbell_st):

the vertex is (-3, 9)